OpenStudy (anonymous):

what number is 7% of 88?

OpenStudy (kira_yamato):

7/100 * 88 = 6.16

OpenStudy (hoblos):

0.07*88 = 6.16

OpenStudy (amistre64):

88/100 = 1% so you just add 7 of them together

OpenStudy (anonymous):

THANKS

OpenStudy (anonymous):

6.16